# ReGAL: Refactoring Programs to Discover Generalizable Abstractions

### v1

@@ Abstract
While large language models (LLMs) are increasingly being used for program synthesis, they lack the global view needed to develop useful abstractions; they generally predict programs one at a time, often repeating the same functionality. Generating redundant code from scratch is both inefficient and error-prone. To address this, we propose Refactoring for Generalizable Abstraction Learning (ReGAL), a gradient-free method for learning a library of reusable functions via code refactorization, i.e. restructuring code without changing its execution output. ReGAL learns from a small set of existing programs, iteratively verifying and refining its abstractions via execution. We find that the shared function libraries discovered by ReGAL make programs easier to predict across diverse domains. On three datasets (LOGO graphics generation, Date reasoning, and TextCraft, a Minecraft-based text-game), both open-source and proprietary LLMs improve in accuracy when predicting programs with ReGAL functions. For CodeLlama-13B, ReGAL results in absolute accuracy increases of $11.5\%$ on graphics, $26.1\%$ on date understanding, and $8.1\%$ on TextCraft, outperforming GPT-3.5 in two of three domains. Our analysis reveals ReGAL’s abstractions encapsulate frequently-used subroutines as well as environment dynamics. [^1] Code available at https://github.com/esteng/regal_program_learning
Program induction, refactoring, tool discovery
@@ 1 Introduction
An increasing range of tasks can be tackled by using a large language model (LLM) to generate an executable program for a given query; this paradigm has been applied in computer vision (Surís et al., 2023; Gupta et al., 2018; Cho et al., 2023), robotics (Ahn et al., 2022; Singh et al., 2023), tool use (Schick et al., 2023; Lu et al., 2023; Qin et al., 2023), and general reasoning (Lyu et al., 2023). In all these cases, the overall program generation framework is the same: an individual query is given (along with an instructive prompt) to an LLM, which produces a program that, when executed, yields the desired result. Crucially, each program is generated independently (as shown in Fig. 1), with no reference to other queries or programs, and is composed of primitive operations, i.e. the domain language’s built-in operations. This approach has two major and related limitations:
1) Lack of Reusability: Each program is designed as a one-off script to solve a given example but is not reused by other examples. This increases redundancy and can result in unnecessary errors: for two examples requiring a shared subroutine, the model might correctly generate the subroutine in one example and make a mistake in the other. For instance, in Fig. 1 (top) although the “primitive-only” model had previously generated nonagons, it uses the wrong interior angle (40.5). ReGAL’s $\tt{draw\_small\_9gon\text{()}}$) function, on the other hand, executes correctly.
@@
Both limitations can be traced to a lack of global context as the model sees each example separately, so it lacks a mechanism for developing reusable abstractions. This differs greatly from how humans write code: generally, developers might start solving individual tasks with one-off solutions, but quickly begin to develop a library of shared abstractions and code snippets for related problems, thereby reducing the redundancy of their code, promoting efficiency and readability (McConnell, 2004; Downey, 2012). Furthermore, functions can be verified: once we have tested a function, we can rely on it in the future – something that is harder to do for ever-changing one-off code snippets. Such abstraction and verification is only sensible if the code synthesis process takes place over the course of multiple examples. In other words, if presented with a single, one-off task, there is no reason not to write a one-off script.
While abstraction offers numerous benefits, it comes with the risk of over-fitting, where a function tailored to a specific example loses its generalizability. For instance, a function like $\tt{draw\_9gon\_snowflake\text{()}}$) may perfectly match one example but fails to generalize. Conversely, $\tt{draw\_small\_9gon\text{()}}$) is a more versatile function applicable in various contexts. The ability to produce novel programs using primitive operations needs to be balanced with the benefits of encoding subroutines into reusable abstractions (O’Donnell, 2015). A similar balance between flexibility and efficiency appears in a variety of domains, including language (O’Donnell, 2015; Yang, 2016), biology (Futuyma & Moreno, 1988), manufacturing (Flynn & Jacobs, 1987), and programming (Ellis et al., 2021). To strike this balance in LLM-based program synthesis, we propose Refactoring for Generalizable Abstraction Learning (ReGAL). ReGAL refines abstractions iteratively by refactoring programs as well as verifying, correcting, and pruning abstractions such that overly specific or incorrect programs are improved upon or removed. ReGAL relies on two key elements: a small set of programs using primitive operations (i.e. primitive programs) and an execution environment (e.g., Python). Importantly, we show ReGAL can learn from LLM-generated programs without requiring any human annotations.
@@
We demonstrate the broad applicability of ReGAL by testing it on three diverse datasets: LOGO (Ellis et al., 2021; Wong et al., 2021), a program induction task; a date reasoning task (Srivastava et al., 2022) known to challenge LLMs (Suzgun et al., 2022); and TextCraft (Prasad et al., 2023), a text-based game for crafting Minecraft objects. Across these tasks, ReGAL significantly improves the accuracy of the predicted programs from various LLMs – especially open-source LLMs – over a baseline that predicts primitive programs (i.e. programs without ReGAL’s abstractions). For instance, CodeLlama-13B’s (Roziere et al., 2023) accuracy increases by $11.5\%$, $26.1\%$, and $8.1\%$ on LOGO, Date, and TextCraft respectively, surpassing larger models like GPT-3.5 (cf. Sec. 5). In Sec. 6, we show that ReGAL’s abstractions are reusable across examples, encapsulate key domain functionalities, and we include an error analysis further highlighting the features making ReGAL effective. Sec. 6.3 reveals that ReGAL can improve over the baseline with minimal training examples, yielding major improvements even with a $50\%$ reduced training set.
@@ 2 Related Work
Program Induction.   Program induction involves learning a symbolic and programmatic mapping of inputs to outputs. Humans are adept at this kind of “rule-learning” (Marcus et al., 1999; Fürnkranz et al., 2012). ReGAL aims to learn a set of general functions that can be used to map inputs to outputs, i.e. a form of program induction. Ellis et al. (2021) present DreamCoder, a method for combining program induction and synthesis that uses a wake-sleep Bayesian learning method to learn programs. Wong et al. (2021) extend this work to incorporate language, using an alignment model as part of the joint model. Like Ellis et al. (2021), Grand et al. (2023) adopt a similar symbolic search procedure, but use an LLM to document abstractions. The symbolic search procedure used by this line of past work has relied on data structures for search that assume the domain language is $\lambda$-calculus (Lake et al., 2015; Ellis et al., 2021; Wong et al., 2021; Grand et al., 2023), which is not typically used for software development. In contrast, ReGAL uses an LLM-based search procedure, allowing us to use more flexible languages like Python, which are more commonly used by developers and better represented in pre-training data.
@@
Program Synthesis and Tool Use.   Tool use by LLMs (Schick et al., 2023; Mialon et al., 2023) refers to a form of program synthesis or semantic parsing where an LLM generates API calls to external tools (e.g. calculators, search functions, etc.). This formulation has also been applied to reasoning tasks (Lyu et al., 2023; Chen et al., 2022) as well as other domains such as computer vision (Surís et al., 2023; Gupta & Kembhavi, 2023; Cho et al., 2023), summarization (Saha et al., 2022), and robotics (Ahn et al., 2022; Singh et al., 2023; Huang et al., 2022, 2023). Past work has attempted to induce tools from examples. Cai et al. (2023) induce tools using an LLM for reasoning tasks from BigBench (Srivastava et al., 2022); unlike our work, their system generates one tool per task. While this can offer benefits for homogenous reasoning tasks (e.g. sorting words alphabetically), heterogenous tasks like the ones we explore require multiple functions. More akin to our work, Yuan et al. (2023) and Qian et al. (2023) induce multiple tools for vision and math tasks using an LLM-based framework which also includes retrieval-based parsing. In addition to focusing on different domains, we place an emphasis on efficiency, highlighting ReGAL’s performance on open-source LLMs (both only consider proprietary models like GPT-3.5). We also differ in our focus on refactoring, and in the amount of information we provide to the refactoring model: unlike Yuan et al. (2023) and Qian et al. (2023), we do not provide in-context examples of the kinds of tools we want the model to create, investigating instead what abstractions the model builds without domain-specific guidance.
@@ 3 Methodology
In this section, we describe the overall pipeline of our method: Refactoring for Generalizable Abstraction Learning (ReGAL). ReGAL consists of two phases: the training or induction stage where abstractions (i.e., helper functions) are learned, and the testing or synthesis stage, where abstractions are used to generate programs for test queries. During training, as illustrated in Fig. 2, ReGAL discovers reusable abstractions by generating candidate helper functions, validating their correctness, and debugging via editing and pruning of ineffective helper functions.
Given a set of demonstrations $(q,p)$ of queries $q$ and gold primitive programs $p$, we first preprocess the data to cluster examples based on query similarity, described in Sec. 3.1. The training stage then builds abstractions by refactoring primitive programs in batches (Sec. 3.2), while the testing stage solves new queries by generating programs that glue together the learned abstractions with primitive operations (Sec. 3.3). We use GPT-3.5 for all our training prompts; at test time we experiment with a range of LLMs, focusing on freely available open-source LLMs.
@@ 3.1 Preprocessing
Clustering Data.   In order to form abstractions that are shared between examples, the refactoring LLM requires a multi-instance scope, i.e. it must receive a batch of related $(q,p)$ tuples at a time. We implement this by clustering examples using an embedding of the query $q$. Specifically, we embed each query using OpenAI’s Ada embedding model (OpenAI, 2022) and hierarchically cluster the embeddings using Ward’s agglomerative clustering algorithm (Ward Jr, 1963), implemented via Scikit-Learn (Pedregosa et al., 2011). This gives us a tree of related examples, which we topologically sort and group into batches of $k$ related examples, where $k$ is a hyperparameter (see Appendix C for all hyperparameter values).
Figure 2: ReGAL starts by refactoring a batch of primitive programs to develop a set of modified programs and helper functions (Stage 1). It then verifies the results of refactored programs, optionally retrying failed programs according to environment feedback. Useful helper functions are added to the Code Bank along with example usage added to the Demo Bank (Stage 2). Periodically, we edit and prune the Code Bank to improve its functions (Stage 3). At test time, ReGAL agent has access to the Code Bank, the Demo Bank, and the remaining original programs. It is compared against a baseline agent which has access to a larger number of original programs.
@@ 3.2 Training
ReGAL’s training data consists pairs of queries $q$ and primitive programs $p$. The training phase outputs: 1) the Code Bank ($C$): the library of helper functions abstracted out during training and 2) the Demo Bank ($D$): examples of the functions being used. As shown in Fig. 2, the training phase is an iterative process where the LLM receives as input a batch of queries and primitive programs and then proposes helper functions that can be abstracted (Stage 1). For each batch, candidate helper functions are evaluated based on the correctness of the refactored programs that occur in (Stage 2a). After verification, failed examples are isolated into a second batch and re-attempted (Stage 2b). To improve the quality of the resultant Code Bank, we periodically edit helper functions after a fixed number of batches to improve their pass rate (over unit tests) and prune ineffective helper functions (Stage 3). At the end of training, the resultant library of helper functions – the code bank $C$ – is stored for further use during testing along with successful demonstrations of refactored programs using helper functions – the demo bank $D$. Note that the training process can be repeated over multiple epochs. Below we describe each stage in detail, with the overall algorithm detailed in Algorithm 1.
@@
Stage (2a): Verification (verify).   To avoid introducing errors, we need to verify the helper functions and refactored programs generated by the LLM by executing them and comparing the results to the original, i.e. determining if $\hat{p}\text{()}=p\text{()}$). The refactored program $(q,\hat{p})$ is stored as a demonstration for future use by the agent (cf. Sec. 3.3) if it passes verification. Only helper functions that pass verification are added to $C$. We also store a record of programs that failed verification, as these will be crucial in $\tt{editCodeBank\text{()}}$) and $\tt{pruneCodeBank\text{()}}$), which improve existing functions and prune functions leading to failures.
Stage (2b): Feedback-based Retrial (retry).   If a program fails to pass verification, we optionally retry the refactoring process. In a follow-up prompt, we present failed programs and their helper functions. We also include environment feedback from execution (i.e. the output or error message produced). [^2] We do not include the output for LOGO as it is an image. The refactoring LLM then produces a new version of each failed program; these are verified and their helpers are added to $C$ if correct.
Stage (3a): Editing Code Bank (editCodeBank).   From the $\tt{verify\text{()}}$) module, some helper functions fail to pass all unit tests because they contain incorrect abstractions; for example, a function like $\tt{draw\_triangle\text{()}}$) might start with a hardcoded value for a small size, leading it to fail on medium triangles. To update such functions, we construct a prompt for each function in $D$ that shows the LLM passing and failing unit tests and asks it to propose edits to the function; this occurs once every $\tt{editEvery}$ iterations, where $\tt{editEvery}$ is a hyperparameter. We replace a function if it passes more unit tests after editing.

### v2

@@ Abstract
While large language models (LLMs) are increasingly being used for program synthesis, they lack the global view needed to develop useful abstractions; they generally predict programs one at a time, often repeating the same functionality. Generating redundant code from scratch is both inefficient and error-prone. To address this, we propose Refactoring for Generalizable Abstraction Learning (ReGAL), a gradient-free method for learning a library of reusable functions via code refactorization, i.e., restructuring code without changing its execution output. ReGAL learns from a small set of existing programs, iteratively verifying and refining its abstractions via execution. We find that the shared function libraries discovered by ReGAL make programs easier to predict across diverse domains. On five datasets – LOGO graphics generation, Date reasoning, TextCraft (a Minecraft-based text-game) MATH, and TabMWP – both open-source and proprietary LLMs improve in accuracy when predicting programs with ReGAL functions. For CodeLlama-13B, ReGAL results in absolute accuracy increases of $11.5\%$ on LOGO, $26.1\%$ on date understanding, and $8.1\%$ on TextCraft, outperforming GPT-3.5 in two of three domains. Our analysis reveals ReGAL’s abstractions encapsulate frequently-used subroutines as well as environment dynamics. [^1] Code: https://github.com/esteng/regal_program_learning.
Program induction, code refactoring, tool discovery, LLMs
@@ 1 Introduction
An increasing range of tasks can be tackled by using a large language model (LLM) to generate an executable program for a given query; this paradigm has been applied in computer vision (Surís et al., 2023; Gupta et al., 2018; Cho et al., 2023), robotics (Ahn et al., 2022; Singh et al., 2023), tool use (Schick et al., 2023; Lu et al., 2023; Qin et al., 2023), and complex reasoning (Lyu et al., 2023). In all these cases, the overall program generation framework is the same: an individual query is given (along with an instructive prompt) to an LLM, which produces a program that, when executed, yields the desired result. Crucially, each program is generated independently (as shown in Fig. 1), with no reference to other queries or programs, and is composed of primitive operations, i.e., the domain language’s built-in operations. This approach has two major and related limitations:
1) Lack of Reusability: Each program is designed as a one-off script to solve a given example but is not reused by other examples. This increases redundancy and can result in unnecessary errors: for two examples requiring a shared subroutine, the model might correctly generate the subroutine in one example and make a mistake in the other. For instance, in Fig. 1 (top) although the “primitive-only” model had previously generated nonagons, it draws a polygon with an incorrect angle. ReGAL’s $\tt{draw\_small\_9gon\text{()}}$) function, on the other hand, executes correctly.
@@
Both limitations can be traced to a lack of global context as the model sees each example separately, so it lacks a mechanism for developing reusable abstractions. This differs greatly from how humans write code: generally, developers might start solving individual tasks with one-off solutions, but quickly begin to develop a library of shared abstractions and code snippets for related problems, thereby reducing redundancy in their code, promoting efficiency and readability (McConnell, 2004; Downey, 2012). Furthermore, functions can be verified: once we have tested a function, we can rely on it in the future – something that is harder to do for ever-changing one-off code snippets. Such abstraction and verification is only sensible if the code synthesis process takes place over the course of multiple examples. In other words, if presented with a single, one-off task, there is no reason not to write a one-off script.
While abstraction offers numerous benefits, it comes with the risk of over-fitting, where a function tailored to a specific example loses its generalizability. For instance, in Fig. 1, a function like $\tt{draw\_9gon\_snowflake\text{()}}$) may perfectly match one example but fails to generalize. Conversely, $\tt{draw\_small\_9gon\text{()}}$) is a more versatile function applicable in various contexts. The ability to produce novel programs using primitive operations needs to be balanced with the benefits of encoding subroutines into reusable abstractions (O’Donnell, 2015). A similar balance between flexibility and efficiency appears in a variety of domains, including language (O’Donnell, 2015; Yang, 2016), biology (Futuyma & Moreno, 1988), manufacturing (Flynn & Jacobs, 1987), and programming (Ellis et al., 2021). To strike this balance in LLM-based program synthesis, we propose Refactoring for Generalizable Abstraction Learning (ReGAL). ReGAL refines abstractions iteratively by refactoring programs as well as verifying, correcting, and pruning abstractions such that overly specific or incorrect programs are improved upon or removed. ReGAL relies on two key elements: a small set of programs using primitive operations (i.e., primitive programs) and an execution environment (e.g., Python). Importantly, we show ReGAL can learn from LLM-generated programs without requiring any human annotations.
@@
We demonstrate the broad applicability of ReGAL by testing it on five diverse datasets: LOGO (Ellis et al., 2021; Wong et al., 2021), a program induction task; a date reasoning task (Srivastava et al., 2022) known to challenge LLMs (Suzgun et al., 2022); TextCraft (Prasad et al., 2023), a text-based game for crafting Minecraft objects; a subset of MATH (Hendrycks et al., 2021) which contains algebra word problems, and TabMWP (Lu et al., 2022), a collection of math quesitons about tables. Across these tasks, ReGAL significantly improves the accuracy of the predicted programs from various LLMs – especially open-source LLMs – over a baseline that predicts primitive programs (i.e., programs without ReGAL’s abstractions). For instance, CodeLlama-13B’s (Roziere et al., 2023) accuracy increases by $11.5\%$, $26.1\%$, and $8.1\%$ on LOGO, Date, and TextCraft respectively, surpassing larger models like GPT-3.5 (cf. Sec. 5). In Sec. 6, we show that ReGAL’s abstractions are reusable across examples, encapsulate key domain functionalities, and we include an error analysis further highlighting the features that make ReGAL effective. Sec. 6.3 reveals that ReGAL can improve over baseline primitive programs with minimal examples, yielding major improvements even with a $50\%$ reduced training set.
@@ 2 Related Work
Program Induction.   Program induction involves learning a symbolic and programmatic mapping of inputs to outputs. Humans are adept at this kind of “rule-learning” (Marcus et al., 1999; Fürnkranz et al., 2012). ReGAL also aims to learn a set of general functions that can be used to map inputs to outputs, i.e., a form of program induction. Ellis et al. (2021) present DreamCoder, a method for combining program induction and synthesis that uses a wake-sleep Bayesian learning method to learn programs. Wong et al. (2021) extend this work to incorporate language, using an alignment model as part of the joint model. Like Ellis et al. (2021), Grand et al. (2023) adopt a similar symbolic search procedure, but use an LLM to document abstractions. The symbolic search procedure used by this line of past work has relied on data structures that assume the domain language is $\lambda$-calculus (Lake et al., 2015; Ellis et al., 2021; Wong et al., 2021; Grand et al., 2023), which is not typically used for software development. In contrast, ReGAL has an LLM-based search procedure, allowing us to use flexible languages like Python, which are more commonly used by developers and better represented in pre-training data.
@@
Program Synthesis and Tool Use.   Tool use by LLMs (Schick et al., 2023; Mialon et al., 2023) refers to a form of program synthesis or semantic parsing where an LLM generates API calls to external tools (e.g., calculators, search functions, etc.). This formulation has also been applied to reasoning tasks (Lyu et al., 2023; Chen et al., 2022) as well as other domains such as computer vision (Surís et al., 2023; Gupta & Kembhavi, 2023; Cho et al., 2023), summarization (Saha et al., 2022), and robotics (Ahn et al., 2022; Singh et al., 2023; Huang et al., 2022, 2023). Past works have attempted to induce tools from examples. Cai et al. (2023) induce tools using an LLM for reasoning tasks from BigBench (Srivastava et al., 2022); unlike our work, their system generates one tool per task. While this can offer benefits for homogenous reasoning tasks (e.g., sorting words alphabetically), heterogenous tasks like the ones we explore require multiple functions. More akin to our work, Yuan et al. (2023), Qian et al. (2023), and Wang et al. (2024) induce multiple tools for vision and math tasks using an LLM-based framework which also includes retrieval-based parsing. In addition to focusing on different domains, we place an emphasis on learning a shared code bank that can be used by multiple LLMs to generate programs including several open-source LLMs. We also differ in our focus on refactoring, and in the amount of information we provide to the refactoring model: unlike Yuan et al. (2023) and Qian et al. (2023), we do not provide in-context examples of the kinds of tools we want the model to create, investigating instead what abstractions the model builds without domain-specific guidance. While Wang et al. (2024) also prunes their learned functions, their method does not involve refactoring; in contrast, we learn functions by refactoring and periodically edit the codebank (in addition to pruning it).
@@ 3 Methodology
In this section, we describe the overall pipeline of our method: Refactoring for Generalizable Abstraction Learning (ReGAL). ReGAL consists of two phases: the training or induction stage where abstractions (i.e., helper functions) are learned, and the testing or synthesis stage, where abstractions are used to generate programs for test queries. During training, as illustrated in Fig. 2, ReGAL discovers reusable abstractions by generating candidate helper functions, validating their correctness, and debugging via editing and pruning of ineffective helper functions. Given a set of demonstrations $(q,p)$ of queries $q$ and gold primitive programs $p$, we first preprocess the data to cluster examples based on query similarity, described in Sec. 3.1. The training stage then builds abstractions by refactoring primitive programs in batches (Sec. 3.2), while the testing stage solves new queries by generating programs that glue together the learned abstractions with primitive operations (Sec. 3.3). We use GPT-3.5 for training; at test time we use a range of LLMs, focusing on freely available open-source LLMs.
@@ 3.1 Preprocessing
Clustering Data.   In order to form abstractions that are shared between examples, the refactoring LLM requires a multi-instance scope, i.e., it must receive a batch of related $(q,p)$ tuples at a time. We implement this by clustering examples using an embedding of the query $q$. Specifically, we embed each query using OpenAI’s Ada embedding model (OpenAI, 2022) and hierarchically cluster the embeddings using Ward’s clustering algorithm (Ward Jr, 1963), implemented via Scikit-Learn (Pedregosa et al., 2011). This gives us a tree of related examples, which we topologically sort and group into batches of $k$, where $k$ is a hyperparameter (see Appendix C for all hyperparameter values).
Figure 2: ReGAL starts by refactoring a batch of primitive programs to develop a set of modified programs and helper functions (Stage 1). It then verifies the results of refactored programs, optionally retrying failed programs according to environment feedback. Useful helper functions are added to the Code Bank along with example usage added to the Demo Bank (Stage 2). Periodically, we edit and prune the Code Bank to improve its functions (Stage 3). At test time, the ReGAL agent has access to the Code Bank, the Demo Bank, and the remaining original programs. It is compared against a baseline agent which has access to a larger number of original programs.
@@ 3.2 Training
ReGAL’s training data consists pairs of queries $q$ and primitive programs $p$. The training phase outputs: 1) the Code Bank ($C$): the library of helper functions abstracted out during training and 2) the Demo Bank ($D$): examples of the functions being used. As shown in Fig. 2, the training phase is an iterative process where the LLM receives as input a batch of queries and primitive programs and then proposes helper functions that can be abstracted (Stage 1). For each batch, candidate helper functions are evaluated based on the correctness of the refactored programs that occur in (Stage 2a). After verification, failed examples are isolated into a second batch and re-attempted (Stage 2b). To improve the quality of the Code Bank, we periodically edit helper functions after a fixed number of batches to improve their pass rate (over unit tests) and prune ineffective helpers (Stage 3). After training, the library of helper functions (Code Bank $C$) is stored for use during testing along with successful demonstrations of programs using helper functions (Demo Bank $D$). Note that the training process can be repeated over multiple epochs. Below we describe each stage in detail, with the overall algorithm detailed in Algorithm 1.
@@
Stage (2a): Verification (verify).   To avoid introducing errors, we need to verify the helper functions and refactored programs generated by the LLM by executing them and comparing the results to the original, i.e., determining if $\hat{p}\text{()}=p\text{()}$). The refactored program $(q,\hat{p})$ is stored as a demonstration for future use by the agent (cf. Sec. 3.3) if it passes verification. Only helper functions that pass verification are added to $C$. We also store a record of programs that failed verification, as these will be crucial in $\tt{editCodeBank\text{()}}$) and $\tt{pruneCodeBank\text{()}}$), which improve existing functions and prune functions leading to failures.
Stage (2b): Feedback-based Retrial (retry).   If a program fails to pass verification, we optionally retry the refactoring process. In a follow-up prompt, we present failed programs and their helper functions. We also include environment feedback from execution (i.e., the output or error message produced). [^2] We do not include the output for LOGO as it is an image. The refactoring LLM then produces a new version of each failed program; these are verified and their helpers are added to $C$ if correct.
Stage (3a): Editing Code Bank (editCodeBank).   From the $\tt{verify\text{()}}$) module, some helper functions fail to pass all unit tests because they contain incorrect abstractions. For example, a function like $\tt{draw\_triangle\text{()}}$) might start with a hardcoded value for a small size, leading it to fail on medium triangles. To update such functions, we construct a prompt for each function in $D$ that shows the LLM passing and failing unit tests and asks it to propose edits to the function; this occurs once every $\tt{editEvery}$ iterations, where $\tt{editEvery}$ is a hyperparameter. We replace a function if it passes more unit tests after editing.